A statistical risk for infectious disease. Please show work and explain  

Calculate the attack rate for person who did eat pizza. Ate Ice Cream Did Not Eat Ice Cream DIARRHEA TOTAL ATE PIZZA DID NOT EAT PIZZA ! DIARRHEA 39 TOTAL 52 15 14 40 30 O 39/52 50/67 O 53/92 O 14/40 O 20/45
Answer #1

1. The rate of diarrhoea who ate pizza and ice cream is 39/52
The rate of diarrhoea who ate pizza only is 14/40
The rate of diarrhoea who ate ice cream only is 11/15
The rate who didn’t eat neither pizza nor ice cream is 9/30
The question is the attack rate for person who ate pizza is the addition of both the attack rate who ate pizza with ice cream and pizza only.
The attack rate is= 39/52 +14/40 = 53/92
The answer is option C.
2. The percentage of rate of diarrhoea who ate pizza and ice cream is 39/52 = 75%
The percentage of rate of diarrhoea who ate pizza only is 14/40 = 35%
The percentage of rate of diarrhoea who ate ice cream only is 11/15 = 73.3%
The percentage of rate of diarrhoea who ate neither pizza nor ice cream is 9/30 = 30%
The answer is option D. Ice cream only because the rate percentage is highest.